OpenStudy (anonymous):

Evaluate: 16 - 2 + 3 - 4

jhonyy9 (jhonyy9):

so you will can do it sure easy just make 16-2 and assum to the result 3 and from this result subtract 4 and will get the right answer good luck bye

OpenStudy (anonymous):

16-2+3-4=16-2-1 =16-3 =13